If sin ​-1 x + sin ​ ​-1 y = π/2 , then ​ x ​2 is equal to 1) y ​2 2)(1-y) ​1/2 3) 1 – y ​2 4) 0

  If sin​-1 x + sin​-1 y = π/2,
  then x​2 is equal to                                                      1) y​2                  2)(1-y)​1/2            3) 1 – y​2     4) 0

Grade:12

2 Answers

Shahid
7 years ago
Dear Tushar , 
We Know that
Sin-1x + cos-1x = \Pi/2
And Given is that Sin-1x + Sin-1y = \Pi/2
 
Therefore , we get that...
Sin-1y = cos-1x
Sin-1y = Sin-1(\sqrt{1 - x^{2}}
 
Therefore
y = \sqrt{1 - x^{2}}
 
On Squaring both sides .
we get
 
x​2  = 1 – y​2
ankit singh
3 years ago
his problem involves both differentiation as well as inverse trigonometric functions.

Now given that,

sin1x+sin1y=π2

sin1x=π2sin1y

sin1x=cos1y

y=cossin1x

dydx=sinsin1x.ddx(sin1x)

dydx=x11x2

dydx=x1x2.